1. Suppose G is an abelian group of order 6 containing an element of order 3. Prove G is cyclic?
  2. Suppose G has only one element (say a) of order 2. show xa=ax for all x in G?
1
Expert's answer
2021-05-31T19:30:47-0400

1. Suppose "G" is an abelian group of order 6 containing an element "a" of order 3. Let us prove "G" is cyclic. By Cauchy's Lemma, the group "G" contains an element "b" of order 2. Since elements "a" and "b" are commute, and 2 and 3 are relatively prime, we conclude that the element "c=ab" is of order "2\cdot 3=6". Therefore, the cyclic subgroup "\langle c\rangle" of "G" has order 6. Since "G" also has order 6, we conclude that "G=\langle c\rangle". Therefore, "G" is cyclic group generated by "c."


2. Suppose "G" has only one element (say "a") of order 2. Let us show "xa=ax" for all "x \in G". Since for each "x\in G", for the element "b_x=xax^{-1}" we have that "b_x^2=b_xb_x=(xax^{-1})(xax^{-1})=xa(x^{-1}x)ax^{-1}=xa^2x^{-1}=xx^{-1}=e", we conclude that for each "x\in G" the element "b_x" has order 2. Since "a" is a unique element of order 2, we conclude that "b_x=a." Therefore, for each "x\in G" we have that "xax^{-1}=a", and hence each "x\in G" we have that "xa=ax".